OpenStudy (anonymous):

y = -2/3x – 2 –2x + 3y = -6 Identify the slope and the y-intercept of the equation.

OpenStudy (ghazi):

to find the slope remember y=mx+c

OpenStudy (ghazi):

so in your first equation slope= coefficient of x...guess what it is?

OpenStudy (anonymous):

umm it would be -2 ? @ghazi

OpenStudy (ghazi):

nope it's -2/3

OpenStudy (anonymous):

okayyy, so what do I do know? @ghazi

OpenStudy (ghazi):

well in the next equation try to write it in standard form that is y= mx+c

OpenStudy (anonymous):

So I would -2x+3y=-6 into standard form? @ghazi

OpenStudy (ghazi):

3y=-6+2x............therefore y= (-6+2x)/3 now what next? solve this

OpenStudy (ghazi):

it will be same as you have done above...and to find intercept try to write equation x/a+y/b=1

OpenStudy (anonymous):

woooah i just got confused..im sorry @ghazi

OpenStudy (ghazi):

divide second equation by -6 your intercept would be -2

OpenStudy (ghazi):

y intercept...=-2